OpenStudy (anonymous):

Solve for x. Enter only a number. please help, will medal for most help thanks loves log 2 + log (2x) = 2

Nnesha (nnesha):

remember \[\log x + \log y = \log (x \times y) \] \[\log x - \log y = \log \frac{ x }{ y }\]

Nnesha (nnesha):

so you have to change log 2 + log (2x) do division or multiplication ??

Nnesha (nnesha):

to*

OpenStudy (anonymous):

log 2 + log 2x = log 2(2x) = log (4x)

OpenStudy (anonymous):

To solve log(4x) = 2, do 10^ on each side: \[10^{\log (4x)} = 10^{2}\] On the left side of the equation, the 10 and log cancel each other out. So this simplifies to 4x = 100.

OpenStudy (anonymous):

so 25